# Structure of Kerr black hole spacetimes in Weyl conformal gravity

### v1

@@ 6 Variation of $\kappa$
Additionally, we append the suffix “a” to a domain to indicate when the cosmological features and regions are no longer present as they have been pushed out beyond $r=+\infty$, making them have an Anti-de Sitter background. For instance, Domain IVa may be thought of as Domain IV but without the cosmological ergosurface $\mathcal{E}_{\mathrm{C}}$ and horizon $\mathrm{H}_{\mathrm{C}}$, and the corresponding cosmological ergoregion $\mathrm{E}$ and cosmological spacelike region $\mathrm{S}^{+}$. This thus gives Domain IVa the structure of black hole spacetimes of GR Kerr-AdS. We present the additional domains appearing with $\kappa<0$ in table 2.
@@
For $\gamma>\frac{2}{3}$. this boundary (28) is, in fact, given by $\gamma_{\mathrm{max}}$, discussed in the previous section. From this, we see that increasing $\kappa>0$ acts to increase $\gamma_{\mathrm{max}}$, while making $\kappa$ more negative decreases $\gamma_{\mathrm{max}}$.
@@
This is of course limited to $\gamma<\frac{2}{3}$ from the restriction in (20). While this represents the upper edge of the Domain V region in the map, we are unfortunately unable to find a similar useful expression for the right edge of Domain V.
@@
Unfortunately, we have also been unable to find a similar closed form expression for the extremal spin limit $\mathrm{H_{\mathrm{ex(s)}}}$, defining the transition between the Domain III and IV spacetimes in figure 4. These may just be solved for on a case by case basis using the procedure outlined in subsection 4.1.
Now that we have described the indicated domains, we may now analyze the effect of increasing $\kappa$ on the spacetime structure. Turning to figure 5, with $(\kappa,a)=(0.01,0.25)$, we see that the increase of $\kappa$ leads to the cosmological features being pulled in to smaller $r$. In comparison to the $\kappa=0$ case in figure 3, we see that even the $\gamma=0$ case has the cosmological horizon $\mathrm{H}_{\mathrm{C}}$ and ergosurface $\mathcal{E}_{\mathrm{C}}$ at finite $r$.
We see that $\gamma_{\mathrm{max}}$ increases above $\gamma_{\mathrm{max}}=1$ when we bring up $\kappa>0$, as $k$ (7) will then switch sign at a higher value of $\gamma$.
@@
When we now turn to $\kappa<0$, we see that domains containing cosmological features are still present. For instance, we still have some Domain IV spacetimes in figure 4, outside the region defining Domain IVa. This is as the quartic de Sitter term may still remain repulsive as long as $k>0$ corresponding to $\kappa>-k_{0}$, with the transition, of course, determined by (28). In general, however, making $\kappa$ more negative acts to push such cosmological features further out, as can be seen in figure 6.
@@ 7 Dependence on the spin $a$
Figure 7 demonstrates the effect of varying spin $a$ on the spacetimes in the parameter space of $\gamma$ and $\kappa$. Obviously, as $\Delta^{\mathcal{E}}$ (8) has no dependence on $a$, we do not see any changes in the presence and ergosurfaces as we increase spin $a$.
@@ 7.1 Comparison to CG Schwarzschild
Next, we see the familiar dashed diagional line given by (25). While in CG Kerr, this boundary is confined to $\gamma<\frac{2}{3}$ due to (20), no such restriction exists for CG Schwarzschild. This boundary thus extends past $\gamma=\frac{2}{3}$ in figure 8. Crossing from below this diagonal line to above it results in the coalescence of the Cauchy horizon of a black hole with its event horizon, leading them to mutually annihilate. Meanwhile, in CG Kerr, for $\gamma<\frac{2}{3}$, (25) marks the annihilation of an outer ergosurface $\mathcal{E}_{\mathrm{O}}$ with a cosmological ergosurface $\mathcal{E}_{\mathrm{C}}$
@@
Most importantly, as we have discussed extensively, while the sign of the quadratic de Sitter term in CG Schwarzschild (3) is only dependent on $\kappa$, the sign of the quartic de Sitter term in CG Kerr (6) is dependent on the full sign of $k$ (7). Thus, the background in CG Kerr is de Sitter when $k>0$, flat when $k=0$, and Anti-de Sitter when $k<0$.
While black hole spacetimes in CG Kerr always have at least two horizons, namely the inner Cauchy horizon $\mathrm{H}_{\mathrm{I}}$ and outer event horizon $\mathrm{H}_{\mathrm{O}}$, the medium blue regions in figure 8 show Schwarzschild black holes with just the event horizon. This difference is of course attributed to the repulsion from the spin $a$ in Kerr spacetimes not being present for Schwarzschild.
The black holes described by the violet domain in figure 8 are of note. These can be thought of as nested black holes. While we have an outer event horizon $\mathrm{H}_{\mathrm{EH,}2}$, a middle horizon $\mathrm{H}_{\mathrm{Middle}}$, there is another innermost event horizon $\mathrm{H}_{\mathrm{EH,}1}$. Thus, the causal structure of this domain would be $\mathrm{S}^{-}\rightarrow\mathrm{T}\rightarrow\mathrm{S}^{-}\rightarrow\mathrm{T}$. While this intermediate horizon $\mathrm{H}_{\mathrm{Middle}}$ demarcates a transition $\mathrm{T}\rightarrow\mathrm{S}^{-}$ as $r$ is increased, and thus is another repulsive horizon, we do not call it a Cauchy horizon. This is because the $\mathrm{T}$ region it encloses does not possess the singularity.
The fact that corollary nested black holes are not found for CG Kerr can again be ascribed to the repulsive spin $a$ dominating the small $r$ behavior, and thus preventing the existence of an innermost $\mathrm{S}^{-}$ region.
In general then, we may say that the differences between the spacetimes present in the $\kappa$ vs $\gamma$ parameter space of CG Schwarzschild and CG Kerr may be owed to three things. Firstly, the fact that ergosurfaces and horizons are distinct in CG Kerr but not CG Schwarzschild. Secondly, we have the ubiquity of repulsion from the presence of spin $a$ in CG Kerr. Finally, we have the fact that the causal structure of the background is determined by the sign of $k$ (7) in CG Kerr, while it is only dependent on the sign of $\kappa$ in CG Schwarzschild.
@@ 8 Conclusions and future prospects
In this work, we elucidated the structure of spacetimes arising in the $\gamma-\kappa-a$ parameter space of the conformal gravity Kerr metric, for positive mass $(\beta>0)$ as evaluated on the equatorial plane $(\theta=\frac{\uppi}{2})$. We found about a dozen distinct configurations of these spacetimes. We further noted that the causal and ergoregion structure of the black hole spacetimes of CG Kerr have corollaries in the Kerr-de Sitter and Kerr-Anti-de Sitter black hole spacetimes of general relativity, depending keenly on the values of $\gamma,\kappa,$ and $a$.
We also derived formulae for the extremal spin limit, extremal horizon cosmological limit, and extremal ergosurface cosmological limit. This allowed us to determine that due to the additional dependence on the $\gamma$ parameter, spacetime solutions of the conformal gravity Kerr metric may possess extremal limit values lying above or below those of corollary general relativity solutions, depending of course on the specific combination of parameters.
Future work could look at expanding our classification of spacetimes to $\beta<0$ and to the $r<0$ region within the ring singularity. Additionally, as in [22], it would be of interest to explore finding the equatorial photon circles within this metric, and how they may further subclassify the spacetimes we found here.
@@ Appendix B Derivation of extremal limit values
Deriving the Empty case for four repeated roots, we consider that all four roots of a general quartic are given by $-c_{3}/(4c_{4})$ when $c_{0}/c_{4}=0$ [60]. This yields

### v2

@@ 6 Variation of $\kappa$
Additionally, we append the suffix “a” to the domain name to indicate when the cosmological features ($\mathrm{H}_{\mathrm{C}})$ and $(\mathcal{E}_{\mathrm{C}}$) and regions ($\mathrm{H}_{\mathrm{C}}$ and $\mathrm{S}^{+}$) are no longer present as they have been pushed out beyond $r=+\infty$, leaving them with an Anti-de Sitter background. For instance, Domain IVa may be thought of as Domain IV but without the cosmological ergosurface $\mathcal{E}_{\mathrm{C}}$ and horizon $\mathrm{H}_{\mathrm{C}}$, and the corresponding cosmological ergoregion $\mathrm{E}$ and cosmological spacelike region $\mathrm{S}^{+}$. This thus gives Domain IVa the structure of black hole spacetimes of GR Kerr-AdS. We present the additional domains appearing with $\kappa<0$ in table 2.
@@
For $\gamma>\frac{2}{3}$, this boundary (40) is, in fact, given by $\gamma_{\mathrm{max}}$, discussed in the previous section. From this, we see that increasing $\kappa>0$ acts to increase $\gamma_{\mathrm{max}}$, while making $\kappa$ more negative decreases $\gamma_{\mathrm{max}}$.
@@
This is of course limited to $\gamma<\frac{2}{3}$ from the restriction in (26). While this represents the upper edge of the Domain V region in the map, we failed to find a similarly useful expression for the right edge of Domain V.
@@
We also failed to find a closed form expression for the extremal spin limit $\mathrm{H_{\mathrm{ex(s)}}}$, defining the transition between the Domain III and IV spacetimes in figure 4. We therefore solved for these using the procedure outlined in subsection 4.1.
Having described the indicated domains, we may now analyze the effect of increasing $\kappa$ on the spacetime structure. Turning to figure 5, with $(\kappa,a)=(0.01,0.25)$, we see that the increase of $\kappa$ leads to the cosmological ergosurfaces $\mathcal{E}_{\mathrm{C}}$ and horizons $\mathrm{H}_{\mathrm{C}}$ being pulled in to smaller $r$. In comparison to the $\kappa=0$ case in figure 3, we see that even the $\gamma=0$ case has the cosmological horizon $\mathrm{H}_{\mathrm{C}}$ and ergosurface $\mathcal{E}_{\mathrm{C}}$ at finite $r$.
Solving for $\gamma_{\mathrm{max}}$ from $k=0$ (12), we see that it increases above $\gamma_{\mathrm{max}}=1$ when we bring up $\kappa>0$. This is as $k$ (12) will then switch sign at a higher value of $\gamma$.
@@
When we now turn to $\kappa<0$, domains containing the cosmological features $\mathrm{H}_{\mathrm{C}}$ and $\mathcal{E}_{\mathrm{C}}$ are still present. For instance, we still have some Domain IV spacetimes in figure 4, outside the region defining Domain IVa. This is as the quartic de Sitter term may still remain repulsive as long as $k>0$ corresponding to $\kappa>-k_{0}$, with the transition, of course, determined by (40). In general, however, making $\kappa$ more negative acts to push $\mathrm{H}_{\mathrm{C}}$ and $\mathcal{E}_{\mathrm{C}}$ further out, as can be seen in figure 6.
@@ 7 Dependence on the spin $a$
Figure 7 demonstrates the effect of varying spin $a$ on the spacetimes in the parameter space of $\gamma$ and $\kappa$. Obviously, as $\Delta^{\mathcal{E}}$ (13) has no dependence on $a$, we see no changes in the presence and ergosurfaces as we increase spin $a$.
@@ 7.1 Comparison to CG Schwarzschild
Next, we see the familiar dashed diagional line given by (31). While in CG Kerr, this boundary is confined to $\gamma<\frac{2}{3}$ due to (26), no such restriction exists for CG Schwarzschild. This boundary thus extends past $\gamma=\frac{2}{3}$ in figure 8. Crossing from below this diagonal line to above it results in the coalescence of the Cauchy horizon of a black hole with its event horizon, leading them to mutually annihilate. Meanwhile, in CG Kerr, for $\gamma<\frac{2}{3}$, (31) marks the annihilation of an outer ergosurface $\mathcal{E}_{\mathrm{O}}$ with a cosmological ergosurface $\mathcal{E}_{\mathrm{C}}$ .
@@
Most importantly, as we have discussed extensively, while the sign of the quadratic de Sitter term in CG Schwarzschild (8) depends only on $\kappa$, the sign of the quartic de Sitter term in CG Kerr (11) depends on the sign of $k$ (12). Thus, the background in CG Kerr is de Sitter when $k>0$, flat when $k=0$, and Anti-de Sitter when $k<0$.
While black hole spacetimes in CG Kerr always have at least two horizons, namely the inner Cauchy horizon $\mathrm{H}_{\mathrm{I}}$ and outer event horizon $\mathrm{H}_{\mathrm{O}}$, the medium blue regions in figure 8 show Schwarzschild black holes with just the event horizon. This difference is of course attributed to the repulsion from the spin $a$ in CG Kerr spacetimes being absent for CG Schwarzschild.
The black holes described by the violet domain in figure 8 are of note. These can be thought of as nested black holes. While we have an outer event horizon $\mathrm{H}_{\mathrm{EH,}2}$, a middle horizon $\mathrm{H}_{\mathrm{Middle}}$, there is another innermost event horizon $\mathrm{H}_{\mathrm{EH,}1}$. Thus, the causal structure of this domain is $\mathrm{S}^{-}\rightarrow\mathrm{T}\rightarrow\mathrm{S}^{-}\rightarrow\mathrm{T}$. While this intermediate horizon $\mathrm{H}_{\mathrm{Middle}}$ demarcates a transition $\mathrm{T}\rightarrow\mathrm{S}^{-}$ as $r$ is increased, and thus is another repulsive horizon, we do not call it a Cauchy horizon because the $\mathrm{T}$ region it encloses does not possess the singularity. The absence of nested black holes in CG Kerr can again be ascribed to the repulsive spin $a$ dominating the small $r$ behavior, and thus preventing the existence of an innermost $\mathrm{S}^{-}$ region.
In general, then, the differences between the spacetimes present in the $\kappa$ vs. $\gamma$ parameter space of CG Schwarzschild and CG Kerr may be attributed to three things. Firstly, ergosurfaces and horizons are distinct in CG Kerr but are coincident in CG Schwarzschild. Secondly, we have the ubiquity of repulsion from the spin $a$ in CG Kerr. Finally, the causal structure of the background is determined by the sign of $k$ (12) in CG Kerr, while it is only dependent on the sign of $\kappa$ in CG Schwarzschild.
@@ 8 Conclusions and future prospects
In this work, we elucidated the structure of spacetimes arising in the $\gamma-\kappa-a$ parameter space of the conformal gravity Kerr metric, for positive mass $(\beta>0)$ as evaluated on the equatorial plane $(\theta=\frac{\uppi}{2})$. We found thirteen distinct configurations of these spacetimes. We further noted that the causal and ergoregion structure of the black hole spacetimes of CG Kerr have analogues in the Kerr-de Sitter and Kerr-Anti-de Sitter black hole spacetimes of general relativity, depending keenly on the values of $\gamma,\kappa,$ and $a$.
We also derived formulae for the extremal spin limit, extremal horizon cosmological limit, and extremal ergosurface cosmological limit. This allowed us to determine that due to the additional dependence on the $\gamma$ parameter, spacetime solutions of the conformal gravity Kerr metric may possess extremal limit values lying above or below those of the analogous general relativity solutions, depending of course on the specific combination of parameters.
In addition to this, we calculated the surface gravities and temperatures associated with Hawking radiation for the horizons in the CG Kerr spacetimes. We found that these quantities both vanish at the extremal horizon limits, as one would also see in general relativity.
Future work could look at expanding our classification of spacetimes to $\beta<0$ and to the $r<0$ region within the ring singularity. As has been done for the CG Schwarzschild metric [27], it may be interesting to see if an appropriate conformal transformation can render curvature scalars at the CG Kerr ring singularity regular.
Additionally, as in [27], it would be of interest to explore finding the equatorial photon circles within this metric, and how they may further subclassify the spacetimes we found here. Studying the efficiency of energy extraction via the Penrose process would also be worthwhile to consider for this metric compared to the corresponding metrics of GR.
@@ Acknowledgments
We also thank Philip D. Mannheim for some instructive insights into conformal gravity.
@@ Appendix B Derivation of extremal limit values
Deriving the Empty case for four repeated roots, we consider that all four roots of a general quartic are given by $c_{3}/(4c_{4})$ when $c_{0}/c_{4}=0$ [77]. This yields